Apple-Space, Inc. year-end balance sheets are as follows:

At December 31                                                2012                       2011                       2010

Assets

Cash                                                                      33,458                   41,104                   41,570

A/R, Net                                                               96,971                   71,234                   54,318

Merchandise Inventory                                 125,655                 90,440                   58,433

Prepaid Expenses                                            11,104                   10,685                   4,665

Plant Assets, net                                              311,468                 285,378                 260,914

Total Assets                                                       578,656                 498,841                 419,900

Liabilities and Equity

Accounts Payable                                             148,408                 85,990                   54,873

Long-term notes payable secured by

Mortgages on plant assets                           106,612                 114,733                 92,798

Common Stock, $10 per value                    162,500                 162,500                 162,500

Retained Earnings                                            161,136                 135,618                 109,729

Total Liabilities and equity                           578,656                 498,841                 419,900                

Apple-Space Inc. income statements for the years ended December 31, 2012 and 2011, follow

For Year Ended December 31                      2012                       2011

Sales                                                                      752,253                 593,621

COGS                                                                    458,874                 385,854

Other Operating Expenses                           233,198                 150,186

Interest Expense                                              12,788                   13,653

Income Taxes                                                    9,779                     8,904    

Total Costs and Expenses                             714,639                 558,597                                

Net Income                                                        37,614                   35,024

Earnings/Share                                                 2.31                        2.16       

Compute the following:

  1. Debt and equity ratios

Debt Ratio

Choose Numerator                         /              Choose Denominator                     =             Debt Ratio %

2012       ___________________                         __________________                                   _________

2011       ___________________                          ___________________                                 _________

                                                                                                                Equity Ratio

Choose Numerator                         /              Choose Denominator                     =             Debt Ratio %

2012       ___________________                       __________________                                   _________

2011       ___________________ ___________________                                 _________

Ref: 9-11
Answer #1

1. Debt ratio = Total debt/ Total assets

2012 :

total debt = 255020(148408+106612)

total assest = 578656 (given)

hence, debt ratio = 255020/578656*100= 44.071%(approx)

2011 :

total debt = 200723(85990+114733)

total asset = 498841

hence, Debt ratio = 200723/498841*100=40.24%(approx)

2. Equity ratio = Total shareholder's equity/ Total assets

2012 :

Total equity = 323636 (578656-255020)

Total assets = 578656

hence, Equity ratio = 323636/578656 = 55.929%(approx)

2011 :

total equity = 298118 (498841-200723)

total assets = 498841

hence, Equity ratio = 298118/498841 = 59.76%(approx)

note:

  1. equity can be also find out by adding common stock and retained earnings
  2. if the question is asked for what is equal to debt then take the equation of debt ratio and equity ratio and change the sides to attain the desired results

the equation thus formed will be :

  • Debt ratio * total assets = debt
  • (1-equity ratio)*total asset = debt

(1- equity ratio) is calculted as the debt ratio and equity ratio both added will attain 100% or 1

since there are only two components which in the liablities side they are debt and equity